OpenStudy (anonymous):

which is the appropriate symbol to be put in the blank make /9/ + /-5/__ /9-5/ true? a.= b.> c.< d.+ ???

OpenStudy (hba):

@danyaennim Welcome to Open Study :)

OpenStudy (anonymous):

thankyou! :)

OpenStudy (anonymous):

HINT: +(-x)=-x

OpenStudy (anonymous):

Does it help

OpenStudy (anonymous):

@sauravshakya i will try it

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

14>4 is this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

But what about your first question

OpenStudy (anonymous):

what question?

OpenStudy (anonymous):

ahhh ok

OpenStudy (anonymous):

i answered it already someone help me

OpenStudy (anonymous):

9+(-5) _____ 9-5

OpenStudy (anonymous):

So, what did u get?

OpenStudy (anonymous):

ahhhhhhhhhkkk

OpenStudy (anonymous):

14>4

OpenStudy (anonymous):

> <=== is the answer

OpenStudy (anonymous):

9+(-5) = 9-5

OpenStudy (anonymous):

9+(-5)= 14 9-5=4 the bigger one is 14 so my answer is greater than (>)

OpenStudy (anonymous):

actually danyaennim is right as the question is in terms of absolute brackets